OpenStudy (anonymous):

Why did "d" orbitals start at "3d" instead of "4d"?

OpenStudy (moongazer):

I think it is because of their energy level

OpenStudy (anonymous):

I mean before 3d is 4s so why not 4d?

OpenStudy (anonymous):

electrons are filled in increasing order of energies, in accordance to the aufbau principle. According to it the filling of electron is governed by "n+l" rule and "n" rule. for 3d orbital, n+l = 3+2 = 5 for 4s orbital, n+l = 4+0 = 4 since the n+l value of 4s is less than that of 3d , 4s is filled first. See the atttachment for order of filling electrons.